Group Selection Page
HELP
Controls Group
Data Analysis Group
ITDM
WP76
DET
Vacuum
EEE
WP78
WP74
FXE
SCS
SPB
MID
HED
SQS
SXP
Sample Environment
Operation
PSPO
XO
TS
Migrated2Zulip
|
Technical integration and coordination
SASE1 Status and SASE3 Readiness
|
Technical integration and coordination logbook, Contact: A. Violante, G. Wellenreuther |
Not logged in |
 |
|
Message ID: 97
Entry time: 10 Feb 2020, 10:12
|
Author: |
Frederik Wolff-Fabris |
Location: |
General |
Severity: |
Information |
Subject: |
Notes Dispatch meeting Feb 10th, 2020 |
|
|
SPB: Nothing to report;
FXE: Everything ongoing; nothing special to be reported;
MID: Finalize motors of split and delay line; MID has to test interlock for vacuum loop update.
HED: Commissioning of laser heat table. It may need assistance from EEE and CTRL during the week.
SCS: Nothing to report - Last week's mounted setup will be disassembled in order to mount the standard setup for users;
EEE: OCD was called from XRO during beamline alignment of SA2: 1 BMPI motor had a problem and is under investigation - EEE reminds on the new homing position for
negative direction motion.
CTRL: Following reset of train-ID and timing system for today.
XPD: Shifted Beam Pointing Studies from A. Koch to Tuesday.
DET: Nothing to report
*machine is currently not running due to filter clean in the cold compressor. Shall be re-started this afternoon and by tomorrow must be lasing.
*Master time reset and train-ID reset this morning. Then EEE (check systems), followed by CTRL (karabo servers re-start) and ITDM (check DAQ operation). LAS and DET are
informed.
*Wednesday: Network stress test with Cameras by CTRL on Wednesday 12:00-14:30. |